OpenStudy (anonymous):

How do I do this? Write the equation of the line that passes through (2, 4) and (1, -3) in slope-intercept form.

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Np :)